OpenStudy (anonymous):

solve cos (theta/2)-1=0

OpenStudy (anonymous):

let me use x instead of theta so we have cos(x/2) -1=0 or cos(x/2)=1 cos(x/2)=cos 0 hence x/2 = 2n(pi) where n is an integer x= 4n(pi) where n is an integer

OpenStudy (cggurumanjunath):